The sum of two positive number is 14 and difference between their squares is 56. What is the sum of their squares?


1
Expert's answer
2021-05-31T05:29:01-0400

Let the numbers be a and b,


a + b = 14     

a2 – b2 = 56    


 (a2 – b2) = (a + b) (a – b)

The difference between numbers  squares is 56.

 (a + b) (a – b) = 56

The sum of two positive number is 14,

so  

(a + b) = 14, (1)

(a – b) = 56 /(a+b)= 56/14= 4     (2)

From formula (2) we have:

a=4+b

From formula (1) we have:

(4 + b + b) = 14

2b=14 - 4=10

b=5,

a=4 + b=4+5=9,


a = 9,  b = 5


Sum of squares =a2 + b2 = 92 + 52 = 81 + 25 = 106 


Answer:

106